Use the method of Laplace transforms and the accompanying proof results to solve the initial value problem. + y +9y + 18y

Af(t) 3- 5 20 10 15 Square wave
Answer #1

noy if f(t+1) = f(t); fet) is function nok with Pri od 7 L ( f(t) = f (t) est de ng T 27 st - fltl et du I f(t) est dt + f f(hlf(t)) = 8W -(0-1BT e Jeck). est de =) 2 n-- (n-1ST stores j f(t) est dt. E ne) 7 S feljest do -54 I Given D.E y + qyt syL (fit) a So, I fit), est at 1 e los Sa estan (570) (1-é5) cité55) E = gust X S (l-e -SS ) lipess) 1 3(less) X (bess) (itess)let K A B St3 Sad sis+15+18) S A ( ² +95 +18) +B(2+651+ c(²+3) s(32+95 + 18) (A+B+C) } + GA +6B +3C) + 18 A s(+95 +18) =) A =So, Y(c) = 3. [ Ž + ج ۱۱-) 2 S+3 stro) 2 a) YCS) = 1 & Chests (3 sh + )] st3 sto L (uct-c). fct-c)) els L (fit)) 2 tl lets. 2